The total volume required to reach the endpoint of a titration required more than the 50 mL total volume of the buret. An initial volume of 49.56 ± 0.05 mL was delivered, the buret was refilled, and an additional 1.61 ± 0.05 mL was delivered before the endpoint was reached. The titration of a blank solution without the analyte required 0.47 ± 0.04 mL. Calculate the endpoint volume corrected for the blank and its absolute uncertainty.

**The answer is not 51.70 +/- 0.09 so please do not state as so
Answer #1

To do this,

set up the equation as follows:

[49.56(± 0.05) + 1.61(±0.05)] - 0.47(± 0.04)

Next:

49.56 + 1.61= 51.7 -0.47= 50.7 ml (3 sig figs)

Finally:

Square root of (0.05)^2 +(0.05)^2 + (0.04)^2 = 0.08 ml (one sig fig)

Hence,

Answer=50.7±0.08ml